Calculate solubility of sold caf2


1
Expert's answer
2021-01-25T04:31:02-0500

Ksp(CaF2) = 4.0 × 10−11


Solution:

CaF2(s) ⇌ Ca2+(aq) + 2F(aq)

The Ksp expression for CaF2(s) is:

Ksp = [Ca2+][F]2 = 4.0×10−11

ICE Table:



Substitute the equilibrium concentration values from the table into the Ksp expression: 

[Ca2+][F]2 = [ x ] × [ 2x ]2 = 4x3 = Ksp

Substitute in the Ksp value to solve for x, the molar solubiltity. 

4x3 = 4.0×10−11

x3 = 1.0×10−11

x = 2.15×10−4 M


Answer: The molar solubiltity of CaF2(s) is 2.15×10−4 M.